OpenStudy (anonymous):

Solve the equation. |2x - 1| = 7

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

and

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

x= 4 and -3

OpenStudy (anonymous):

add 1 to both sides the -1 cancels out so you have 2x=8 divide both sides by 2 the 2 cancels out and 8/2=4

OpenStudy (anonymous):

basically yes solve it like 10 said but also |2x - 1| = -7

OpenStudy (anonymous):

so -3 and 4 am i correct

OpenStudy (anonymous):

2x-1=7 and 2x-1=-7

OpenStudy (anonymous):

Yup -3 and 4 is the answer

OpenStudy (anonymous):

Thnxs to all the pple that helped i want to give medals to all of yall

OpenStudy (anonymous):

It dont matter about the medal as long as you understood :)

OpenStudy (anonymous):

thnxs

OpenStudy (anonymous):

:)